OpenStudy (grandmoffchris):

1. -2r = 4 2. 4z = 12 3. 2b = 10 4. 6p = -72 5. 2/3m = -8 6. a/16 = -20

OpenStudy (grandmoffchris):

@pooja195 Pls help

OpenStudy (grandmoffchris):

@girlmeetsworld @imqwerty @mittens23309

pooja195 (pooja195):

1) divide both sides by -2

OpenStudy (grandmoffchris):

it would be -2?

pooja195 (pooja195):

Yes r=-2 Do the same for all

OpenStudy (grandmoffchris):

Oh

OpenStudy (grandmoffchris):

For all of em?

pooja195 (pooja195):

2. 4z = 12 divide both sides by 4 z=? 3. 2b = 10 divide both sides by 2 b=? 4. 6p = -72 Divide both sides by 6 p=? 5. 2/3m = -8 Divide both sides by 2/3 m=? 6. a/16 = -20 Multiply by the reciprocal a=-16 x -20

OpenStudy (grandmoffchris):

Number 2 and 3 i made but

pooja195 (pooja195):

but..?

OpenStudy (grandmoffchris):

Nvm

pooja195 (pooja195):

6/6=p -72/6=?

OpenStudy (grandmoffchris):

-72/6 = 1 right?

OpenStudy (grandmoffchris):

no wait

pooja195 (pooja195):

dividing...

OpenStudy (grandmoffchris):

-12

pooja195 (pooja195):

good so 4) p=-12

OpenStudy (grandmoffchris):

Thx

pooja195 (pooja195):

5) Multiply by the reciprocal 

pooja195 (pooja195):

m=?

OpenStudy (grandmoffchris):

mmhhhh

OpenStudy (grandmoffchris):

-6.5

pooja195 (pooja195):

no...

OpenStudy (grandmoffchris):

Oh

pooja195 (pooja195):

\[\huge~\rm~ \frac{ (3)(-8) }{ (2)(1)}\]

pooja195 (pooja195):

-8 x 3=?

OpenStudy (grandmoffchris):

Ok, im not good at math, specially fractions and dividing, just always try but cant.

OpenStudy (grandmoffchris):

and im 7th grade

OpenStudy (grandmoffchris):

fractions involve dividing, dividing idk how to divide

pooja195 (pooja195):

-8 x 3=?

OpenStudy (grandmoffchris):

-24

pooja195 (pooja195):

2 x 1 = 2 -24/2=?

OpenStudy (grandmoffchris):

-12

pooja195 (pooja195):

Good so m=-12

OpenStudy (grandmoffchris):

can you please teach me how to divide?

pooja195 (pooja195):

After we're done check this out http://www.mathsisfun.com/numbers/fractions-division-whole-numbers.html

OpenStudy (grandmoffchris):

Oh thanks :D

pooja195 (pooja195):

6. a/16 = -20 reciprocal equals 16/a a/16 x 16/a=a -20(16)=?

OpenStudy (grandmoffchris):

what? sorry but im kinda

OpenStudy (grandmoffchris):

confused

OpenStudy (grandmoffchris):

i rlly am

OpenStudy (grandmoffchris):

halp?

pooja195 (pooja195):

-20 x 16= ?

OpenStudy (grandmoffchris):

-320?

pooja195 (pooja195):

good so a=-320

OpenStudy (grandmoffchris):

Ok

pooja195 (pooja195):

Thats all ;p

OpenStudy (grandmoffchris):

Wait what

OpenStudy (grandmoffchris):

oh yeah :P

pooja195 (pooja195):

:P

OpenStudy (grandmoffchris):

7. -2x = -15 8. 6x = 42 9. g/3 = -0.7 10. y/4 = 3 11. -7 = j/2 12. y/-2 = -6

OpenStudy (grandmoffchris):

@pooja195 @imqwerty

OpenStudy (grandmoffchris):

help?

pooja195 (pooja195):

@hwyl would gladly help u ;)

OpenStudy (grandmoffchris):

@hwyl

pooja195 (pooja195):

@Watermelon_Talk

OpenStudy (grandmoffchris):

help?

pooja195 (pooja195):

Try them by yourself show your work and show what you get

OpenStudy (grandmoffchris):

Ok

OpenStudy (grandmoffchris):

4 is 6 - 12 = -72

pooja195 (pooja195):

4?

OpenStudy (grandmoffchris):

3 is 2 x 5 = 10 and 2 is 4 x 3 = 12

OpenStudy (grandmoffchris):

im confused :/

pooja195 (pooja195):

we are on number 7

OpenStudy (grandmoffchris):

i told you im not that good at this stuff,

OpenStudy (grandmoffchris):

im not good at dividing or cractions

OpenStudy (grandmoffchris):

fractions*

pooja195 (pooja195):

7. -2x = -15 When isolating variables like this you need to DIVIDE both sides by -2 -2x = -15 ======== -2 = -2 -2/-2= ?

OpenStudy (grandmoffchris):

-2 / -2 = 1?

pooja195 (pooja195):

Yes good job that leaves us with just x x=-15/-2 a negative / negative can be made into a postivie so x=15/2

OpenStudy (grandmoffchris):

this is for 7?

pooja195 (pooja195):

Yes

OpenStudy (grandmoffchris):

so its x = 15 / 2?

pooja195 (pooja195):

yes

OpenStudy (grandmoffchris):

can u help me with 9? ill do 8,

OpenStudy (grandmoffchris):

i think its just multiplying

pooja195 (pooja195):

ok first tell me what u get for 8

OpenStudy (grandmoffchris):

7 is

OpenStudy (grandmoffchris):

6 x 7 = 42

OpenStudy (grandmoffchris):

@dan815 is the guy that can help me

OpenStudy (grandmoffchris):

he is a human calculator

pooja195 (pooja195):

Well if you dont want my help its fine with me.

OpenStudy (grandmoffchris):

oh yes i do

OpenStudy (grandmoffchris):

need ur help @pooja195 :(

OpenStudy (grandmoffchris):

didnt mean to hurt u

pooja195 (pooja195):